OpenStudy (anonymous):

Which value is a solution for the equation tan x/2 = -1

OpenStudy (anonymous):

tanx is equal to 1 when x = 3pi/4 or -pi/4

OpenStudy (anonymous):

So in your case x would equal 3pi/2 and -pi/2

OpenStudy (anonymous):

because you're dividing by 2 as well.

OpenStudy (anonymous):

oops, I mean tanx is equal to -1 when x is equal to 3pi/4 and -pi/4.

OpenStudy (anonymous):

If tan a = N, one value of a is arctan(a), and we can generalize a= k(pi) + arctan(a) where k is an interger. So here x/2 = k(pi) + arctan(-1) x/2 = k(pi) -45 x= 2kpi-90

OpenStudy (anonymous):

what about this one?

OpenStudy (anonymous):

sorry my computer is being stupid today.

OpenStudy (anonymous):

The value \[\frac{ \Pi }{ 4 }\] is a solution for the equation \[3\sqrt{2}\cos \theta+2=-1\].

OpenStudy (anonymous):

True or false??

OpenStudy (anonymous):

\[\cos(\pi/4) = \frac{ \sqrt{2} }{ 2 }\] \[3\sqrt{2}\frac{ \sqrt{2} }{ 2 } + 2 = 3+2 = 5\] False

OpenStudy (anonymous):

\[\frac{ \sqrt{2} }{ 2 } * \sqrt{2} = 1; fyi\]

OpenStudy (anonymous):

that being said is there a solution for sec x = 0??

OpenStudy (anonymous):

No

OpenStudy (anonymous):

It ranges from 1 to infinity and -infinity to -1. Nothing in between.

OpenStudy (anonymous):

so it would be no solution for the problem?

OpenStudy (anonymous):

Yes.